# A compact model of Escherichia coli core and biosynthetic metabolism

### v2

@@ Abstract
Metabolic models condense biochemical knowledge about organisms in a structured and standardised way. As large-scale network reconstructions are readily available for many organisms of interest, genome-scale models are being widely used among modellers and engineers. However, these large models can be difficult to analyse and visualise, and occasionally generate hard-to-interpret or even biologically unrealistic predictions. Out of the thousands of enzymatic reactions in a typical bacterial metabolism, only a few hundred comprise the metabolic pathways essential to produce energy carriers and biosynthetic precursors. These pathways carry relatively high flux, are central to maintaining and reproducing the cell, and provide precursors and energy to engineered metabolic pathways. Here, focusing on these central metabolic subsystems, we present a manually-curated medium-scale model of energy and biosynthesis metabolism for the well-studied prokaryote Escherichia coli K-12 MG1655. The model is a sub-network of the most recent genome-scale reconstruction, iML1515, and comes with an updated layer of database annotations, as well as a range of metabolic maps for visualisation. We enriched the stoichiometric network with extensive biological information and quantitative data, enhancing the scope and applicability of the model. In addition, here we assess the properties of this model in relation to its genome-scale parent and demonstrate the use of the network and supporting data in various scenarios, including enzyme-constrained flux balance analysis, elementary flux mode analysis, and thermodynamic analysis. Overall, we believe this model holds the potential to become a reference medium-scale metabolic model for E. coli.
@@
Abbreviations: EFM: Elementary Flux Mode; FBA: Flux Balance Analysis; GEM: Genome-scale models; MDF: Max-Min Driving Force; RMSE: Root Mean Squared Error; 3GP: glycerol 3-phosphate; 3hmrsACP: (R)-3-Hydroxytetradecanoyl-ACP; 3MOB: 3-Methyl-2-oxobutanoate; Ala: L-alanine; Asp: L-aspartate; ATP: adenosine triphosphate; F6P: fructose 6-phosphate; G3P: glyceraldehyde 3-phosphate; HdeACP: cis-hexadec-9-enoyl-ACP; palmACP: palmitoyl-ACP; PEP: phosphoenolpyruvate; Ru5P: D-ribulose 5-phosphate.
@@ 1 Introduction
Metabolic models are a valuable tool for biologists and biotechnologists aiming to elucidate and engineer cell metabolism [sarkar_engineering_2019, gu_current_2019, lu_silico_2023]. In their simplest form, such models encode basic biochemical knowledge, such as network structure, reaction stoichiometries, or known reaction directionalities, in a structured and standardised format. However, the scope of these models is often wider, including information about catalysing enzymes and kinetic parameters, as well as annotations that link model elements to external databases. The fast development of high-throughput experimental and computational pipelines has led to genome-scale reconstructions of metabolism now existing for a wide range of microorganisms of interest [gu_current_2019, fang_reconstructing_2020]. One of them, Escherichia coli, has been the most studied prokaryotic organism and, as such, its metabolism has been the subject of extensive modelling efforts spanning over three decades [reed_thirteen_2003, feist_genomescale_2007, orth_comprehensive_2011, monk_iml1515_2017]. In particular, for the common laboratory strain E. coli K-12 MG1655, the most recent genome-scale reconstruction, iML1515, accounts for 2712 enzyme-catalysed reactions mapped in detail to 1515 genes [monk_iml1515_2017].
Genome-scale metabolic network models (GEMs) provide a comprehensive picture of cell metabolism, and constraint-based modelling algorithms using these models have shown remarkable predictive power, for example, when predicting gene essentiality [bernstein_evaluating_2023]. Nevertheless, working with such large models comes with some disadvantages. In the absence of sufficient constraining or parametrisation, simulations based on large networks can easily lead to biologically unrealistic solutions, which may confound the interpretation of modelling results. For example, when designing and testing gene knockout strategies, genome-scale networks often wrongly predict unphysiological metabolic bypasses which have to be manually inspected and filtered out. Another issue is that, owing to their size and complexity, analysis of genome-scale networks is often limited to relatively simple modelling frameworks, such as flux balance analysis (FBA), that can only answer a limited range of questions. More complex methods, including kinetic modelling, sampling of metabolic flux distributions, or elementary flux mode (EFM) analysis [wortel_metabolic_2018], can be used to gain additional insight into the governing principles and constraints of microbial metabolism, but are difficult to apply to large models. Finally, genome-scale models are often hard to visualise comprehensively, which can make the interpretation of computed flux distributions cumbersome and unintuitive.
@@
Here, we introduce iCH360 (named, as per convention, by the initials of the authors followed by the number of genes covered by the model), a manually curated, “Goldilocks-sized” model of E. coli K-12 MG1655 energy and biosynthesis metabolism. The model was derived from the most recent genome-scale reconstruction (iML1515 [monk_iml1515_2017]) and includes all pathways required for energy production and the biosynthesis of the main biomass building blocks, such as amino acids, nucleotides, and fatty acids, but accounts only implicitly for the conversion of these precursors into more complex biomass components. We extended the coverage of annotations pointing to external databases from iML1515, and built custom metabolic maps to facilitate visualising the model and its subsystems [jahan_development_2016]. We complemented the stoichiometric network structure with a curated layer of biological information about catalytic function, protein complex composition, and small-molecule regulation. Finally, we enriched the model with useful quantitative data, including thermodynamic and kinetic constants. Thanks to all these extra layers of information, our model can support a wide range of modelling efforts beyond simple stoichiometric ones.
Below we present the model and demonstrate several use-cases across various modelling scenarios, including enzyme-allocation predictions, EFM analysis, and thermodynamic analysis. The model is freely available in SBML, JSON, and SBTab standard formats, and can be directly used with popular metabolic modelling tools such as the COBRA toolbox [ebrahim_cobrapy_2013].
@@ 2 Results
2.1 A compact model of E. coli energy and biosynthesis metabolism
To assemble the iCH360 model, we started from the core metabolic reactions present in ECC and extended them with the pathways required for the biosynthesis of the main biomass building blocks, including the twenty amino acids, the four nucleotides, and both saturated and unsaturated fatty acids (Fig. 1, Table 1, and Supplementary Figures 1-4). In addition, while not performing a comprehensive review of iML1515, we applied a small number of corrections to the original reactions based on literature knowledge (Supplementary Information A.1).
@@
Figure 1: Metabolic map of the iCH360 model, built with the metabolic visualisation tool Escher [king_escher_2015] and showing the metabolic subsystems included in the model. Shaded areas denote metabolic subsystems already present in the ECC model [orth_reconstruction_2010]. Reaction and metabolite names were omitted from the plot for clarity. Overlaid onto the map is a flux distribution for aerobic growth on glucose computed via parsimonious FBA.
@@
Amino acids biosynthesis Biosynthesis of all 20 amino acids from core metabolism precursors Supplementary Figure 1
Nucleotide biosynthesis Biosynthesis of purine and pyrimidine nucleotides (and deoxynucleotides) from core and amino acid metabolism Supplementary Figure 2
Fatty-acids biosynthesis Biosynthesis of saturated and unsaturated fatty acids present in E. coli from acetyl-CoA Supplementary Figure 3
C1 metabolism One-carbon metabolism Supplementary Figure 4
The final assembled model (Figure 1) comprises $304$ compartment-specific metabolites and $323$ metabolic reactions mapped to $360$ genes, thus qualifying as a medium-scale model ranging in between ECC and iML1515 (Supplementary Figure 5). While similar in scale, our model and ECC2 present a fundamental structural difference. ECC2 was built by systematically removing reactions from its genome-scale parent (iJO1366) [hadicke_ecolicore2_2017]. Thus, its metabolic space spans the production of all compounds consumed in the iJO1336 biomass reaction. In contrast, the metabolic space of iCH360 only reaches biomass building blocks, without explicitly considering peripheral pathways, such as assembly of cell-membrane components, de novo synthesis of cofactors, and active transport of ions in the cell.
To make the model comparable to its parent model iML1515, we computed an equivalent biomass reaction, wherein all biomass requirements not present in our model are converted into an equivalent metabolic cost in terms of precursors, based on manually curated bioproduction routes (Table 2, Supplementary Information A.2).
@@ 2.2 Range of metabolic conversions described by production envelopes
To check how its significantly smaller size and complexity affects the solution space of our model, we first looked at the maximally achievable biomass production flux under a range of growth conditions (Supplementary Figure 6). Across most conditions considered, the model can achieve biomass fluxes comparable to its genome-scale parent. The main differences exist in anaerobic growth on fumarate, alpha-ketoglutarate (AKG) and glycerol, where our model predicts zero growth, while the GEM achieves some (albeit small) biomass production rate. In practice, fermentation in these scenarios is in fact biologically unrealistic (e.g. [boecker_enabling_2022]). The reduced model is thus a good basis for metabolic simulation frameworks with few constraints such as FBA, since the reduction procedure, at least in this test, limits the original solution space of the GEM to more physiologically relevant regions.
To investigate the metabolic capabilities of our model even further, we looked at production envelopes (projections of the model’s solution space onto a smaller set of dimensions, also known as phenotypic phase planes), describing production rates for biomass and a range of metabolites. Figure 2 shows the computed envelopes for aerobic growth on glucose. The reduced model has a production envelope comparable with iML1515, where the main apparent difference is in the production of acetate. These differences can be ascribed to the absence of a series of degradation pathways in our model. These pathways allow the genome-scale model to channel additional carbon flux towards acetyl-CoA and, as a result, to reach higher theoretical acetate production yields of up to $\approx 3$ mole acetate per mole of glucose. Since the physiological relevance of such high acetate yields, in the absence of heterologously expressed pathways, has already been questioned [hadicke_ecolicore2_2017], we chose not to include in the model those additional reactions that would allow for higher acetate yields. A similar pattern of agreement with iML1515 was found when comparing production envelopes under different growth conditions (Supplementary Figure 7).
@@
Figure 2: The iCH360 model shows similar but more precise metabolic capabilities than iML1515. Considering glucose as feedstock and studying ethanol, lactate, and succinate production, a production envelope analysis yields similar results in the two models (note that the dashed line representing the production envelope of iML1515 is sometimes hidden behind the coloured lines). Growth rate and production fluxes were computed by bounding the glucose uptake rate to a maximum of 10 mmol/gDW/h. In the scenario of acetate production (top right panel) iCH360 does not yield an unrealistically high production flux [hadicke_ecolicore2_2017] as predicted iML1515.
@@ 2.3 Connecting reactions to their catalysing enzymes and the enzymes’ protein components
Metabolic models often contain annotations that connect model elements to entries in biological databases, such as EcoCyc [keseler_ecocyc_2017], KEGG [kanehisa_kegg_2017] and MetaNetX [moretti_metanetxmnxref_2021]. However, even for the subset of reactions included in iCH360, the annotations present in iML1515 were incomplete and in part outdated. To fill these gaps, we extended and corrected the original annotations through a mixture of automated querying and manual curation (Figure 3A). Notably, the annotations to EcoCyc, the main knowledgebase for E. coli, are complete so that every metabolic reaction in iCH360 is uniquely mapped to a corresponding EcoCyc ID (with the exception of three reactions, for which a match on the database could only be found for a different choice of redox cofactor, namely NADH instead of NADPH). Additionally, we found $134$ deprecated annotations pointing to the MetaNetX namespace, which were consequently updated with the most up-to-date IDs.
@@
Figure 3: Layers of annotation and biological knowledge supporting the stoichiometric model in iCH360. A: Annotations for the model reactions point to the EcoCyc, metaNetX, and KEGG databases. Bars show the numbers of annotations, highlighting the share of annotations that were added to or updated from the parent model iML1515. B: Some of the biological knowledge parsed from EcoCyc (and manually curated) included in the model-supporting functional annotation graph. The graph captures catalytic relationships between reactions and enzymes, protein subunit compositions, protein-gene mappings, and small-molecule regulation interactions, among others. Shown here as an example are the branches of the graph corresponding to the Glutamate Dehydrogenase (GLUDy) and Glutamate Synthase (GLUSy) reactions. C: Examples of catalytic relationships functionally annotated as either primary or secondary in the graph. D: Functional annotation of catalytic edges as primary or secondary supports phenotypic predictions. Left: Classification of catalytic edge disruptions in the network resulting from simulated knockout of genes associated with essential reactions in the model across $9$ growth conditions (see main text for a description of each disruption class). Right: comparison of predicted disruption outcomes against a large dataset of mutant fitness data [price_mutant_2018] shows that significantly different fitness changes are associated with the different types of disruption in the network.
@@
Leveraging the complete mapping with the EcoCyc database, we parsed, assembled, and manually curated a data structure that enhances the stoichiometric model with a detailed layer of information about enzymes, polypeptides, and genes related to the network (Methods). This data structure takes the form of a weighted graph, wherein nodes represent biological entities (reactions, proteins, genes or compounds) and edges represent (potentially quantitative) functional relationships between them (Methods, Figure 3B, Supplementary Tables 1 and 2), such as catalysis, regulation, protein modification, and protein subunit composition. This graph contains the gathered information in a unified form, allowing users to perform a number of tasks related to metabolic modelling applications. For example, by explicitly mapping reactions to their catalysing enzymes rather than to the associated genes (which may additionally include protein activators or cofactors), it simplifies the definition of meaningful enzyme capacity constraints in the model. Similarly, since the graph topology implicitly defines associations between reactions and genes, boolean gene-protein-reaction (GPR) rules needed for in silico knockout studies can be generated accounting for both catalytic and non-catalytic requirements for each reaction (Supplementary Information A.3). Crucially, these GPR rules can be re-generated as needed whenever the graph is updated with new nodes or edges. Finally, we employed it to estimate the abundance of protein complexes included in the model from measured polypeptide abundances (Supplementary Information A.4) using an automated procedure (see Result section 2.5).
2.4 Primary and secondary catalysis
@@
Through this annotation graph, $318$ metabolic reactions in the model are linked to $289$ catalysing enzymes, with more than $25\%$ of the reactions being catalysed by multiple enzymes (isozymes). Such enzymatic redundancy plays an important role, for example, when designing metabolic engineering strategies to prevent flux through a pathway. However, the different isoenzymes of a reaction need not all be equally important, and treating them as completely equivalent may generate inaccuracies in some phenotypic predictions. For example, while phosphofructokinase activity in E. coli (reaction PFK) is known to be carried by two isozymes, pfkA and pfkB, the latter is known to be responsible only for minor activity under normal physiological conditions [babul_phosphofructokinases_1978]. If these differences are not taken into account, metabolic models can overpredict redundancy in the network, for instance when assessing phenotypic predictions in response to gene knockouts [monk_iml1515_2017].
@@
To address this issue and make the model usable in a wider range of modelling scenarios, we qualified catalytic edges in the graph as either primary or secondary catalysis (Methods, Figure 3C). A catalytic relationship between a reaction and an enzyme was annotated as secondary whenever the enzyme, according to experiments, accounts only for negligible activity for the reaction in the wild-type strain. As experimental evidence we considered in vitro and in vivo complementation studies. Through this curation process, a total of $72$ catalytic edges were functionally annotated as secondary.
@@
To test how this functional annotation can support phenotypic predictions, we investigated the disruption of each essential reaction in the model resulting from the knockout of each one of its associated genes (Methods). We classified the outcome of each possible knockout as: (i) complete disruption, if a reaction loses all of its catalytic edges, (ii) full primary disruption, if a reaction loses all of its primary catalysis edges, but secondary ones are left, (iii) partial primary disruption, if the reaction loses some, but not all, of its primary catalysis edges, or (iv) secondary disruption, if some or all secondary edges are disrupted, but none of the primary ones. We classified the effect of in silico knockouts for aerobic growth on $9$ different carbon sources and compared the results against a large dataset of mutant fitness data, obtained via competitive fitness assays, from [price_mutant_2018].

### v3

@@ Abstract
Metabolic models condense biochemical knowledge about organisms in a structured and standardised way. As large-scale network reconstructions are readily available for many organisms, genome-scale models are being widely used among modellers and engineers. However, these large models can be difficult to analyse and visualise and occasionally generate predictions that are hard to interpret or even biologically unrealistic. Of the thousands of enzymatic reactions in a typical bacterial metabolism, only a few hundred form the metabolic pathways essential to produce energy carriers and biosynthetic precursors. These pathways carry relatively high flux, are central to maintaining and reproducing the cell, and provide precursors and energy to engineered metabolic pathways. Focusing on these central metabolic subsystems, we present iCH360, a manually curated medium-scale model of energy and biosynthesis metabolism for the well-studied bacterium Escherichia coli K-12 MG1655. The model is a sub-network of the most recent genome-scale reconstruction, iML1515, and comes with an updated layer of database annotations and with a range of metabolic maps for visualisation. We enriched the stoichiometric network with extensive biological information and quantitative data, enhancing the scope and applicability of the model. In addition, we assess the properties of this model in comparison to its genome-scale parent and demonstrate the use of the network and supporting data in various scenarios, including enzyme-constrained flux balance analysis, elementary flux mode analysis, and thermodynamic analysis. Overall, we believe this model holds the potential to become a reference medium-scale metabolic model for E. coli.
@@
Abbreviations: EFM: Elementary Flux Mode; FBA: Flux Balance Analysis; GEM: Genome-scale models; MDF: Max-Min Driving Force; PMO: Probabilistic Metabolic Optimisation; RMSE: Root Mean Squared Error; 3GP: glycerol 3-phosphate; 3hmrsACP: (R)-3-Hydroxytetradecanoyl-ACP; 3MOB: 3-Methyl-2-oxobutanoate; Ala: L-alanine; Asp: L-aspartate; ATP: adenosine triphosphate; F6P: fructose 6-phosphate; G3P: glyceraldehyde 3-phosphate; HdeACP: cis-hexadec-9-enoyl-ACP; palmACP: palmitoyl-ACP; PEP: phosphoenolpyruvate; Ru5P: D-ribulose 5-phosphate.
@@ 1 Introduction
Metabolic models are a valuable tool for biologists and biotechnologists who want to elucidate and engineer cell metabolism [sarkar_engineering_2019, gu_current_2019, lu_silico_2023]. In their simplest form, such models encode basic biochemical knowledge, such as network structure, reaction stoichiometries, or known reaction directionalities, in a structured and standardised format. However, the scope of these models is often wider, including information on catalysing enzymes and kinetic parameters, as well as annotations that link model elements to external databases. Rapid development of high-throughput experimental and computational pipelines has led to genome-scale metabolic network reconstructions now existing for a wide range of microorganisms [gu_current_2019, fang_reconstructing_2020]. One of them, Escherichia coli, has been the most studied prokaryotic organism and, as such, its metabolism has been the subject of extensive modelling efforts spanning over three decades [reed_thirteen_2003, feist_genomescale_2007, orth_comprehensive_2011, monk_iml1515_2017]. In particular, for the common laboratory strain E. coli K-12 MG1655, the most recent genome-scale reconstruction, iML1515, accounts for 2712 enzyme-catalysed reactions mapped in detail to 1515 genes [monk_iml1515_2017].
Genome-scale metabolic network models (GEMs) provide a comprehensive picture of cell metabolism, and constraint-based modelling algorithms that use these models have shown remarkable predictive power, for example when predicting gene essentiality in bacteria [bernstein_evaluating_2023]. However, working with such large models comes with some disadvantages. In the absence of sufficient constraining or parameterisation, simulations based on large networks can easily lead to biologically unrealistic solutions. For example, when designing and testing gene knockout strategies, genome-scale networks often wrongly predict unphysiological metabolic bypasses that must be manually inspected and filtered out [he_optimized_2020, satanowski_awakening_2020] (see Supplementary Table 1 for some examples). Another issue is that, owing to their size and complexity, analysis of genome-scale networks is often limited to relatively simple modelling frameworks, such as flux balance analysis (FBA), that can only answer a limited range of questions. More complex methods, including sampling of metabolic flux distributions, elementary flux mode (EFM) analysis [wortel_metabolic_2018], or kinetic modelling can be used to gain additional insight into the governing principles and constraints of microbial metabolism, but are difficult to apply to large models. Finally, genome-scale models are often hard to visualise comprehensively, which can make the interpretation of computed flux distributions cumbersome and unintuitive.
@@
Here, we introduce iCH360 (named, according to convention, by the initials of the authors followed by the number of genes covered by the model), a manually curated “Goldilocks-sized” model of E. coli K-12 MG1655 energy and biosynthesis metabolism. The model was derived from the most recent genome-scale reconstruction (iML1515 [monk_iml1515_2017]) and includes all pathways required for energy production and for the biosynthesis of the main biomass building blocks, such as amino acids, nucleotides, and fatty acids, while the conversion of these precursors into more complex biomass components is described by an effective biomass-producing reaction. We extended the coverage of annotations that point to external databases from iML1515, and built custom metabolic maps to facilitate visualisation of the model and its subsystems [jahan_development_2016]. We complemented the stoichiometric network structure with a curated layer of biological information on catalytic function, protein complex composition, and small molecule regulation. Finally, we enriched the model with useful quantitative data, including thermodynamic and kinetic constants. Thanks to all these extra layers of information, our model can support a wide range of modelling methods beyond simple stoichiometric ones.
In the following, we present the model and demonstrate several use cases across various modelling scenarios, including enzyme-allocation predictions, EFM analysis, and thermodynamic analysis. The model is freely available in the standard formats SBML and JSON and can be used directly with popular metabolic modelling tools such as the COBRA toolbox [ebrahim_cobrapy_2013].
@@ 2 Results
2.1 A compact model of Escherichia coli energy and biosynthesis metabolism
To assemble the iCH360 model, we started from the core metabolic reactions present in ECC and extended them with a curated set of pathways required for the biosynthesis of the main biomass building blocks, including the twenty amino acids, the four nucleotides, and both saturated and unsaturated fatty acids (Fig. 1, Table 1, and Supplementary Figures S1-S4). On the other hand, we specifically omitted from our model the pathways required for the biosynthesis of complex biomass components and polymers, most degradation pathways, the pathways involved in the de novo biosynthesis of cofactors, and the reactions involved in the uptake of metals and ions. In addition, while not performing a comprehensive review of iML1515, we applied a small number of corrections to the original reactions based on the knowledge from the literature (Supplementary Information A.1).
@@ 2.1 A compact model of Escherichia coli energy and biosynthesis metabolism
Figure 1: Metabolic map of the iCH360 model, built with the metabolic visualisation tool Escher [king_escher_2015] and showing the metabolic subsystems included in the model. Shaded areas denote metabolic subsystems already present in the ECC model [orth_reconstruction_2010]. Reaction and metabolite names were omitted from the plot for clarity. Overlaid onto the map is a flux distribution for aerobic growth on glucose, computed via parsimonious FBA.
@@
Amino acids biosynthesis Biosynthesis of all 20 amino acids from core metabolism precursors Supplementary Figure S1
Nucleotide biosynthesis Biosynthesis of purine and pyrimidine nucleotides (and deoxynucleotides) from core and amino acid metabolism Supplementary Figure S2
Fatty-acids biosynthesis Biosynthesis of saturated and unsaturated fatty acids present in E. coli from acetyl-CoA Supplementary Figure S3
C1 metabolism One-carbon metabolism Supplementary Figure S4
The final assembled model (Figure 1) comprises $304$ compartment-specific metabolites ($254$ chemically unique compounds) and $323$ metabolic reactions mapped to $360$ genes, thus qualifying as a medium-scale model ranging in between ECC and iML1515 (Supplementary Figure S5). Although similar in scale, our model and ECC2 present a fundamental structural difference. ECC2 was built by systematically removing reactions from its genome-scale parent (iJO1366) [hadicke_ecolicore2_2017]. Thus, its metabolic space spans the production of all compounds consumed in the iJO1336 biomass reaction. In contrast, the metabolic space of iCH360 only reaches biomass building blocks, without explicitly considering peripheral pathways such as the assembly of cell-membrane components, de novo synthesis of cofactors, and active transport of ions in the cell. To make the model comparable to its parent model iML1515, we constructed an equivalent biomass reaction, in which all biomass requirements not present in our model are summarised by an equivalent metabolic cost in terms of precursors, based on manually curated bioproduction routes (Table 2, Supplementary Information A.2, and Supplementary File S1).
@@ 2.2 Range of metabolic conversions described by production envelopes
To check how its significantly smaller size and complexity affect the solution space of our model, we first looked at the maximum achievable biomass production flux under a range of growth conditions (Supplementary Figure S6). Across most conditions considered, the model achieves biomass fluxes comparable to those of its genome-scale parent. The main differences exist in anaerobic growth on fumarate, alpha-ketoglutarate (AKG), and glycerol, where our model predicts zero growth, while the GEM achieves some (albeit small) biomass production rate. In practice, fermentation in these scenarios is biologically unrealistic (e.g. [boecker_enabling_2022]). The reduced model is thus a good basis for metabolic simulation frameworks with few constraints such as FBA, since the reduction procedure, at least in this test, limits the original solution space of the GEM to more physiologically relevant regions.
To further investigate the metabolic capabilities of our model, we looked at production envelopes (projections of the model solution space onto a smaller set of dimensions, also known as phenotypic phase planes) describing production rates for biomass and a range of metabolites (Methods). Figure 2 shows the resulting envelopes for aerobic growth on glucose. The reduced model has production envelopes comparable to iML1515, except for the production of acetate, where the genome-scale model can achieve considerably higher yields, both aerobically and anaerobically. In order to understand the cause of these differences, we investigated optimal acetate production routes in both models using Flux Balance Analysis. Under aerobic conditions, we found that the differences can be traced to different production abilities for acetyl-CoA, a precursor of acetate. In iCH360, acetyl-CoA is produced entirely from pyruvate oxidation via either the Pyruvate Dehydrogenase (PDH) or the Pyruvate-Formate-Lyase (PFL) reactions (Supplementary Figure S7A, left). On the other hand, iML1515 can produce acetyl-Coa via a number of additional pathways not included in our model (Supplementary Figure S7A, right). These include: the degradation of threonine either directly to acetyl-CoA (THRD, GLYAT), or indirectly via acetaldehyde (THRA, THRD, ATHRDHr, THRA2), which is then converted to acetyl-CoA; the degradation of 2-deoxy-D-ribose 5-phosphate into acetaldehyde (DRPA), followed by conversion into acetyl-CoA; the degradation of autoinducer 2 into acetyl-CoA (AI2K, PAI2I, PAI2T). Indeed, we found that, in the region of the production envelope where the two models diverge significantly, up to $50\%$ of acetyl-CoA production in the genome-scale model is accounted for by these degradation pathways. This is in disagreement with the understanding of the existing literature that under aerobic growth on glucose, most of the acetyl-CoA is produced from oxidation of pyruvate [krivoruchko_microbial_2015, zhang_metabolic_2019]. Confirming these findings, the simultaneous deletion of four reactions blocking these degradation pathways (DRPA, GLYAT, PAI2T, THRA, THRD) in iML1515 brings aerobic acetate production to levels comparable to iCH360 (Supplementary Figure S7C).
In the anaerobic scenario, we found that the differences between the two models are exacerbated by the ability of the genome-scale model to produce more pyruvate (which in turn results in higher acetyl-CoA production) than iCH360. Investigating the cause of this difference, we found that, under anaerobic conditions, the iML1515 solution involves the uptake of external CO2 and its use as a sink for electrons produced in glycolysis (Supplementary Figure S7C), which is thermodynamically unrealistic under ambient CO2 levels. Blocking CO2 uptake reduces the maximal anaerobic pyruvate yield in the genome-scale model, but does not fully close the gap with the production capabilities of our model (Supplementary Figure S7D), implying the existence of additional routes for anaerobic production of pyruvate that are not included in iCH360. A similar pattern was found when comparing production envelopes between the two models under different growth conditions (Supplementary Figure S8).
Overall, the higher maximal yields achievable by iML1515 for acetyl-CoA (aerobically) and pyruvate (anaerobically) appear unrealistic. Hence, in agreement with what was previously done with ECC2 [hadicke_ecolicore2_2017], we chose not to include in iCH360 any additional reactions that would allow for higher acetate yields.
Production envelopes generated with iCH360 were also comparable, albeit not identical, with those computed on other existing medium-scale models, namely ECC and ECC2 (Supplementary Figures S9). Particularly, the maximal yields achievable for each product were nearly identical for the products considered, both aerobically and anaerobically. Some differences between the envelopes can be observed for solutions with higher predicted growth rates. However, since the three models are equipped with different biomass reactions sourced from different parental models (and hence predict different biomass yields), these differences are expected.
@@
Figure 2: The iCH360 model shows similar, but more precise metabolic capabilities than iML1515. Considering glucose as a feedstock and studying ethanol, lactate, and succinate production, a production envelope analysis yields similar results in the two models (note that the dashed line representing the production envelope of iML1515 is sometimes hidden behind the coloured lines). Growth rate and production fluxes were computed by limiting the glucose uptake rate to a maximum of 10 mmol/gDW/h. In the scenario of acetate production (top right panel) iCH360 avoids an unrealistically high production flux [hadicke_ecolicore2_2017] as predicted by iML1515.
@@ 2.3 Connecting reactions to their catalysing enzymes and the enzymes’ protein components
Metabolic models often contain annotations that connect model elements to entries in biological databases, such as EcoCyc [keseler_ecocyc_2017], KEGG [kanehisa_kegg_2017] and MetaNetX [moretti_metanetxmnxref_2021]. However, even for the subset of reactions included in iCH360, the annotations present in iML1515 were incomplete and, in part, outdated. To fill these gaps, we extended and corrected the original annotations through a mixture of automated querying and manual curation (Figure 3A). Notably, the annotations pointing to EcoCyc, the main knowledgebase for E. coli, are complete: every metabolic reaction in iCH360 is uniquely mapped to a corresponding EcoCyc ID (with the exception of three reactions, for which a match on the database could only be found for a different choice of redox cofactor, namely NADH instead of NADPH). Additionally, we found $134$ deprecated annotations pointing to the MetaNetX namespace, which were consequently updated with the most up-to-date IDs.
@@
Figure 3: Layers of annotation and biological knowledge supporting the stoichiometric model in iCH360. A: Annotations for the model reactions point to the EcoCyc, metaNetX, and KEGG databases. Bars show the numbers of annotations, highlighting the share of annotations that were added to or updated from the parent model iML1515. B: Some of the biological knowledge parsed from EcoCyc (and manually curated) included in the model-supporting functional annotation graph. The graph captures catalytic relationships between reactions and enzymes, protein subunit compositions, protein-gene mappings, and small-molecule regulation interactions, among others. Shown here as an example are the branches of the graph corresponding to the Glutamate Dehydrogenase (GLUDy) and Glutamate Synthase (GLUSy) reactions. C: Examples of catalytic relationships functionally annotated as either primary or secondary in the graph. Note that all catalytic relationships were classified as primary by default, unless sufficient evidence was found to annotate them as secondary. D: Functional annotation of catalytic edges as primary or secondary supports can be used to improve phenotypic predictions. Left: Classification of catalytic edge disruptions in the network resulting from simulated knockout of genes associated with essential reactions in the model across $9$ growth conditions (see text for a description of each disruption class). Right: comparison of predicted disruption outcomes against a large data set of mutant fitness data [price_mutant_2018] shows that the different types of disruption tend to lead to significantly different fitness changes.
@@
Using the complete mapping with the EcoCyc database, we parsed, assembled, and manually curated a knowledge graph that enhances the stoichiometric model with a detailed layer of information about enzymes, polypeptides, and genes related to the network (Methods). This data structure takes the form of a weighted graph, where nodes represent biological entities (reactions, proteins, genes, or compounds) and edges represent (potentially quantitative) functional relationships between them (Methods, Figure 3B, Supplementary Tables 2 and 3), such as catalysis, regulation, protein modification, and protein subunit composition. This graph contains the information collected in a unified form, allowing users to perform a number of tasks that occur in metabolic modelling applications. For example, by explicitly mapping reactions to their catalysing enzymes rather than to the associated genes (which may additionally include protein activators or cofactors), it simplifies the definition of meaningful enzyme capacity constraints in the model. Similarly, since the graph topology implicitly defines associations between reactions and genes, boolean gene-protein-reaction (GPR) rules needed for in silico knockout studies can be generated that account for catalytic and noncatalytic requirements for each reaction (Supplementary Information A.3). Crucially, these GPR rules can be regenerated as needed whenever the graph is updated with new nodes or edges. Finally, we used it to estimate the abundance of protein complexes included in the model from the measured polypeptide abundances using a simple automated procedure (Supplementary Information A.4, Results section 2.4).
@@
Through this annotation graph, $318$ metabolic reactions in the model are linked to $289$ catalysing enzymes, with more than $25\%$ of the reactions being catalysed by multiple enzymes (isozymes). Such enzymatic redundancy plays an important role, for example, when designing metabolic engineering strategies to prevent flux through a pathway. However, the different isoenzymes of a reaction need not all be equally important, and treating them as completely equivalent may generate inaccuracies in some phenotypic predictions. For example, while phosphofructokinase activity in E. coli (reaction PFK) is known to be carried by two isozymes, pfkA and pfkB, the latter is known to be responsible only for minor activity under normal physiological conditions [babul_phosphofructokinases_1978]. If these differences are not taken into account, metabolic models can overpredict redundancy in the network, for example, when predicting phenotypes after gene knockouts [monk_iml1515_2017].
@@
To address this issue and make the model usable in a wider range of modelling scenarios, we classified the catalytic edges in the graph as either primary or secondary catalysis (Methods, Figure 3C). The catalytic relationship between a reaction and an enzyme was annotated as secondary whenever the enzyme, according to experiments, accounts only for negligible activity for the reaction in the wild-type strain. As experimental evidence we considered in vitro and in vivo complementation studies. Through this curation process, a total of $72$ catalytic edges were functionally annotated as secondary.
@@
To test how this functional annotation can support phenotypic predictions, we investigated the disruption of each essential reaction in the model, resulting from the knockout of each of its associated genes (Methods). We classified the outcome of each possible knockout as: (i) complete disruption, if a reaction loses all of its catalytic edges; (ii) full primary disruption, if a reaction loses all of its primary catalysis edges, but secondary ones are left; (iii) partial primary disruption, if the reaction loses some, but not all, of its primary catalysis edges, or (iv) secondary disruption, if some or all secondary edges are disrupted, but none of the primary ones. We classified the effect of in silico knockouts for aerobic growth on $9$ different carbon sources and compared the results against a large data set of mutant fitness data, obtained via competitive fitness assays, from [price_mutant_2018].